OpenStudy (anonymous):

Which measure is greatest? A. 7 kg B. 68.1 cg C. 700 mg D. 6,900 g

OpenStudy (mathstudent55):

One way to compare is to convert all measures to the same unit. Let's convert every measure to grams.

OpenStudy (anonymous):

7kg

OpenStudy (mathstudent55):

The last one is already in grams, so we don't need to touch D.

OpenStudy (mathstudent55):

A. \(7~ kg \times \dfrac{1000~g}{1~kg} = 7000~g\) B. \(68.1 ~cg \times \dfrac{1~g}{100~cg} = 0.681~g\) C. \(700 ~mg \times \dfrac{1~g}{1000~mg} = 0.7~g\) D. \(6,900 g \)

OpenStudy (mathstudent55):

Now that all measures are in grams, it's easy to compare.